OpenStudy (noahbred):

Please help? me Frank’s quarter grades are 82.5, 94.7, and 87.9. What grade, rounded to the nearest percent, will Frank have to earn in the last quarter to receive at least a 89% final course grade? A. 90% B. 91% C. 92% D. 89%

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

1) Add 82.5, 94.7, and 87.9 2) divide them by 3 3) subtract that from 89 4) add that to 89

OpenStudy (mathmale):

1. Multiply both sides of the equation by 4. 2. Add: 82.5+94.7+87.9. 3. Subtract this sum from both sides of the equation and simplify The remainder is the unknown grade, x.

OpenStudy (anonymous):

1) Add 82.5, 94.7, and 87.9 = 265.10 2) divide them by 3 = 88.37 3) subtract that from 89 = 0.63 4) add that to 89 = 89.63 check your work 88.37+89.63 =178 / 2 = 89 Rounding 89.63 to nearest whole number, you find that a grade of 90% is needed to average an 89